Shift Summary, Day Shift 07/16/1999

Anthony Barker (tonyb@fnal.gov)
Fri, 16 Jul 1999 15:48:35 -0500 (CDT)

Shift Summary for Day Shift, 07/16/99

On Shift: Elliott Cheu, Tony Barker

Hours of beam available at all intensities: 4.0
Hours of beam available at nominal intensity: 3.0
Hours of E832 data (special runs not included): 3.8
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Brief HCC problem; failed tape drive.

========================================================================
Runs ended this shift:

Run 13921 - BEAM_832 (4.3M events) tape change.
Run 13922 - BEAM_832 BAD run. HCC not working.
Run 13923 - BEAM_832 Almost no data. Killed due to arrival of
DLT repairperson.

Tapes brought to FCC this shift: none.
========================================================================

Accesses during shift: One, during beam down time.
========================================================================

Overview of shift: 832 data taking at 4.0 E12

Beam was erratic generally.

Adjusted -24V trimpot on RR160-NIM2 from
DVM-measured -24.9V to -23.9V in order to
eliminate out-of-range error from SDAQ.
We note that DVM measurements are about
1.7V smaller than SDAQ reported values
for both the +-24V supplies in this NIM bin.

Blue ice replaced at noon. It melts rapidly
in this hot weather.

Excitement at end of shift:

-- Run aborts due to failure of DLT drive
ktev1/3d1

-- Jason L. reloads Banana maps; changes to
Banana T0 offsets.

-- Following run fails dramatically with
bizarre HCC errors; warnings for high
L2BEAM/L1BEAM; fixed by reloading HCC.

-- DLT repairperson arrives; kill run.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13921 14:20 Run ABORTS! Failure of DLT drive ktev1/3d1

13922 001 15:00 New run starts after reloading Banana maps.
Trigger 4 now has ZERO rate; other triggers
strangely screwed up. Finally fixed by
reloading HCC.

13923 001 15:26 Start new E832_BEAM run: triggers normal now.

13923 003 15:30 Run killed due to arrival of DLT repairperson.

========================================================================
Follow up Notes on Previous Shift

Now only one red button on the SDAQ screen should be ignored
(S9-UCLA-PS4 -- bad HA shaper-splitter supply).

========================================================================
Items to Follow up on Next Shift

Keep an eye on the E/P (csigain) of CsI 3114.
Keep an eye on CsI 3114, 961, and Ring Counter timing. Also watch
the temperture of the alcohol and moderate with blue ice as necessary.

Argon-ethane pressure is down to 60 psi: Dave Northacker has
been informed.